### v2

# Observing hidden neuronal states in experiments

Abstract
In this article, we revisit voltage-clamp electrophysiology, a feedback-control method that has successfully been applied to study neuronal (non)linear current-voltage relationships. Yet its full scope has not been explored to date. We reinterpret the voltage-clamp protocol under slow variation of the feedback reference signal using the language of multiple-timescale dynamical systems theory. Gauged under this new epistemic step, we unveil non-observable (hidden) states. More specifically, the experimental measurements from voltage clamp can be put into correspondence with predictions of both stable and unstable states from dynamical systems theory. We demonstrate a general protocol for constructing systematically experimental steady-state bifurcation diagrams for electrophysiologically active cells. We perform our experiments on entorhinal cortex neurons, both excitatory (pyramidal neurons) and inhibitiory (interneurons). Moreover, employing a current-clamp open-loop electrophysiology protocol with slow variation of an applied electrical current drives the neuron to dynamically transition between its observable rest states and its observable spiking states, hence revealing experimental dynamic bifurcations. Overlaying the two aforementioned protocols produces a map of all these neuronal states and effectively leads to an experimental validation of the classical slow-fast dissection method introduced by J. Rinzel in the 1980s and routinely applied ever since in order to analyse slow-fast neuronal models. Our approach opens doors to observing further complex hidden states with more advanced control strategies, allowing to control real cells beyond pharmacological manipulations.
Introduction.–
When characterising the dynamics of nonlinear systems, one fundamental criterion for a model is if its stable states such as stationary solutions or periodic orbits match experimental observations. The ability to validate models is, thus, greatly expanded by experimental tools with the capacity to unveil non-observable (sensitive or dynamically unstable) states that are otherwise inaccessible to standard measurements. The combination of observable and non-observable states gives access to an experimental equivalent of parameter-dependent families of stable and unstable states in a model, which are usually referred to as a bifurcation diagram.
This article applies a new experimental technique of using feedback control to find unstable states to electrophysiology experiments on neuronal cells. Our aim is to support systematic validation of neuron models by comparing bifurcation diagrams and observing their between-cell variability. We focus on unstable parts of input-dependent families of steady-state solutions obtained by feedback-controlled experiments and compare them with indirect evidence from standard measurements from open-loop experiments. This extends recent work of Ori et al. [1, 2] constructing phase diagrams from neuronal data, and complements other approaches such as using data to verify the bifurcation structure of neuronal models [3, 4], model-based data analysis [5] or parameter estimation from data [6].
Our technique is simplification of the so-called control-based continuation (CBC) method, an approach which has been recently demonstrated in mechanical experiments [7], vibrations and buckling experiments [8], pedestrian flow experiments [9], atomic-force microscopy [10], cylindrical pipe flow simulations [11], and feasibility studies for synthetic gene networks [12, 13]. Indeed, CBC is a procedure that combines feedback control [14] and pseudo-arclength continuation [15] in a model-free environment, that is, only reliant upon noisy experimental data in a closed-loop control setup. The objective is to compute experimental bifurcation diagrams, that is, both families of stable and unstable states (either stationary or periodic) together with bifurcation points joining such families. Importantly, the control is noninvasive in the sense that it vanishes once an equilibrium or periodic of the uncontrolled system has been reached; the control is decreased iteratively using a Newton’s method. In our case, the approach is simpler and does not require the use of Newton iterations since we consider a slow ramp on the control target. Hence we exploit de presence of two timescales in the experimental procedure in order to obtain directly an approximation of the steady-state bifurcation curve, and we explain why such a bifurcation curve is obtained at low cost; see below.
In our electrophysiology experiments on entorhinal cortex neurons we treat the cell as an electric circuit, and apply a voltage clamp (VC) [16, 17], followed by current clamp (CC). In the VC setup the electrode acts as a voltage source at the neural membrane, fixing the potential across the neural membrane, measuring the current, while the CC setup adds a fixed external current, measuring the resulting membrane potential; see Fig. 1(a) and the Supplementary Information (SI) for further details. In-vivo neurons are subject to current signals that drive spiking (oscillatory) or rest (steady) states of the neural membrane potential [18]. This makes the CC setup the open-loop part of the protocol. In contrast, the VC part of our experiment the closed-loop feedback-controlled part of the protocol because the voltage source regulates the external current to maintain the hold voltage. VC has been applied successfully to study neuronal nonlinear current-voltage relationships, so-called N-shaped $I$-$V$ characteristics, which cause enhanced neuronal excitability and influence the regenerative activation of certain ionic currents (e.g., sodium) [19, 20, 21, 22, 23] across the neural membrane, into and out of the cell. We show that the VC protocol with a slowly varied reference voltage signal gives access to stable and unstable neuronal steady states of the neuron, which was hinted at in [24, 25]. In contrast, the open loop CC protocol with a slowly varying applied current always follows stable (observable) states, driving the neuron to dynamically transition between its observable rest states and its observable spiking states.
We interpret these combined experimental protocols (VC and CC) using multiple-timescale dynamics, in particular, the dissection method [26], which reveals the dynamic bifurcations in the experiment; see Fig. 1(b).
Figure 1: (a) Experimental setup with brain slice (A), patch pipette (B), reference electrode ((Ag-AgCl pellet) connected to ground (C), amplifier (E, Multiclamp 700B) with CV-7B headstage (D), AD-converter (F, National Instruments NI USB-6343) and standard PC computer (G). (b) VC and CC protocol runs for cell $5$: $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$-curve for VC run (thin bright blue: unfiltered data with sampling time step $5\times 10^{-5}$ s, blue/red/brown: median of $I_{\mathrm{vc}}$ over moving windows of size $\Delta w=4\times 10^{3}$ steps equalling $0.2$s) and $(I_{\mathrm{h}}(t),V_{\mathrm{cc}}(t))$-curve for CC run (orange, mean of $I_{\mathrm{h}}$ over moving windows. (c) $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$-curves for VC protocol of all $5$ cells on waterfall $V_{\mathrm{h}}$-axis (same color coding as (b)).
In this slow-fast framework, the states traced by the VC protocol with slow variation correspond to the steady-state experimental bifurcation diagram of the so-called fast subsystem of the mathematical model describing the protocol [26], that is, the model system with constant input current. Hence, the N-shaped I-V relation of a neuron should be seen as a S-shaped V-I bifurcation diagram; see again Fig. 1(b).
Following this strategy, we demonstrate the feasibility of tracking a family of neuronal steady states (stable and unstable) via variations of reference signal and reparameterizing the obtained curve using the feedback current.
Results.–
Figure 1(b) shows the time profiles $(I_{\mathrm{h}}(t),V_{\mathrm{cc}}(t))$ of the CC protocol run (orange, thin) and $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$ of the VC protocol run (bright blue, thin) for cell $5$ overlaid in the $(I,V)$-plane. After smoothing the VC time profile is the S-shaped curve $(I_{\mathrm{vc,sm}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ (blue/brown/red, thick). It equals the $(I,V)$-characteristic of the stationary neuronal states of the CC protocol, including dynamically unstable states (brown and red). The transition to stable spiking states is compatible with type-I excitability, however we do not have sufficient data to distinguish the different types of excitability; see SI, Fig. S2. Their dynamical instability is inferred in two ways: (i) by the negative slope of the $(I_{\mathrm{vc,sm}},V_{\mathrm{h}})$-curve (brown) after smoothing over a moving window with larger size $\Delta w=1.5\times 10^{4}$ steps ($=0.75$s), or (ii) by the presence of oscillations (neuronal firing) in the CC run at $I_{\mathrm{h}}$ equalling the $I_{\mathrm{vc,sm}}$ (red). The stability boundaries of the stationary states are labelled as bifurcations in Fig. 1(b). The change of stability near the disappearance of stable spiking states at $I_{\mathrm{db}}$ is a Hopf bifurcation. The fold points of the $(I_{\mathrm{vc,sm}},V_{\mathrm{h}})$-curve are saddle-node (fold) bifurcations. Figure 1(c) shows the stationary-state curves with their inferred dynamical stability for all $5$ cells (see SI for CC run time profiles used to partially infer stability of cells 11114444). The cells are vertically ordered and numbered according to depth of the S-shape, determining if they fall into the class of type-I or type-II neurons. Figure 1(c) demonstrates wide variability in steady-state curve shape among cells of nominally same function.
Figure 2: (a) Experimental bifurcation diagrams for an interneuron from the entorhinal cortex. (b) Experimental bifurcation diagrams for a type-II PY neuron from the same region. Protocols identical to Fig. 1(b).
What is more, the method described in this work is versatile and rests upon only standard electrophysiological procotols. Hence, it can be applied to any neuron type. Fig. 2(a) below illustrates this point by showing the results of the VC and CC protocols applied to an inhibitory neuron from the entorhinal cortex while panel (b) shows that the same protocols can be applied to a type-II neuron. The steady-state curve is seemingly more complicated than for PY neurons’ recordings shown on other figures in this supplemental material, as well as, in the main manuscript, however it is compatible with a type-I excitable neuron. Noteworthy, there are examples of such scenarios in the literature, in particular with low-threshold spiking neurons [27]. We also have recordings (not shown) from the PY neurons of the dentate gyrus showing a similar outcome once the result of the VC protocol is superimposed onto the result of the CC protocol.
Analysis.–
To see why VC-run time profiles approximate the experimental bifurcation diagram with unstable stationary states of neurons in CC protocol, we use the formalism of multiple-timescale dynamical systems. Superimposing the data from VC and CC protocol also gives the first experimental illustration of slow-fast dissection. The effect of the respective clamps can be understood in a general class of conductance-based models for the neuron,
$C\dot{V}$ $=-\textstyle{\sum}_{j}I_{j}(x_{j},V)+I_{\mathrm{ext}},$ (1)
$\tau_{j}(V)\dot{x}_{j}$ $=x_{\infty,j}(V)-x_{j},$
describing the current balance across the neuron’s membrane. The membrane potential is $V$, $I_{j}(x_{j},V)_{j=1,\ldots,N}$ are the currents across different voltage-gated ion channel types and $I_{\mathrm{ext}}$ is the external current. Each channel type $j$ has a set of associated gating variables $x_{j}$, which are possibly vectors of length $n_{j}$ if the channel gate has both activation and inactivation states, with steady-state gating functions $x_{\infty,j}(V)$ (also of size $n_{j}$) and relaxation times $\tau_{j}(V)$ (a diagonal $n_{j}\times n_{j}$ matrix). The observed dynamic effects such as oscillations (firing/spiking) and negative-slope $(I,V)$ characteristics are determined by these channel coefficients $I_{j}$, $x_{\infty,j}$ and $\tau_{j}$ that are traditionally obtained by parameter fitting from VC experiments, a difficult and ill-posed problem, as gating variables $x_{j}$ are not directly measured [28].
The VC and CC protocols use different mechanisms for generating $I_{\mathrm{ext}}(t)$. The VC protocol is a closed loop where a voltage source regulates $I_{\mathrm{ext}}$ with high gain $g_{\mathrm{c}}$ to achieve the slowly varying hold voltage $V_{\mathrm{h}}$ at the voltage source for (1), measuring $I_{\mathrm{vc}}$:
$\begin{split}I_{\mathrm{ext}}&\approx I_{\mathrm{vc}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V),\\ \dot{V}_{\mathrm{h}}&=\varepsilon\Delta_{V}\end{split}$ (2)
which turns (1), (2) into a multiple-timescale dynamical system with $1+\sum n_{j}$ fast state variables $(V,x_{j})$ (where $\sum n_{j}$ is the overall number of gating variables) and one slow state variable $V_{\mathrm{h}}$, corresponding to the feedback reference signal [29]. The speed at which $V_{\mathrm{h}}$ varies is $\varepsilon\Delta_{V}$ with $\Delta_{V}(t)=0.183$ mV///ms, where we extract the dimensionless small factor $\varepsilon=10^{-2}$.
In contrast, the CC protocol holds $I_{\mathrm{ext}}$, measuring the generated voltage $V_{\mathrm{cc}}$, thus, corresponding to an open-loop system, permitting e.g. the spiking seen in Fig. 1(b):
$\begin{split}I_{\mathrm{ext}}&\approx I_{\mathrm{h}},\\ V_{\mathrm{cc}}&\approx V,\\ \dot{I}_{\mathrm{h}}&=\varepsilon\Delta_{I}.\end{split}$ (3)
The applied hold current $I_{\mathrm{h}}$ is varied slowly at speed $\varepsilon\Delta_{I}$ with $\varepsilon=10^{-2}$ and $\Delta_{I}=0.66\ldots 0.75$ pA///ms. System (1), (3) is also a slow-fast system with $1+N$ fast variables $(V,x_{j})$ and $1$ slow variable $I_{\mathrm{h}}$. The gain $g_{\mathrm{c}}$ in (2) is limited by the imperfect conductance across the non-zero spatial extent of the membrane. Even though (1) is for the potential $V$ across the entire membrane and only $V_{\mathrm{h}}$ at the clamp is measured, we approximate $V_{\mathrm{h}},V_{\mathrm{cc}}\approx V$ for, the membrane potential, and $I_{\mathrm{h}},I_{\mathrm{vc}}\approx I_{\mathrm{ext}}$ for the external current in (1).
Following a classical multiple-timescale approach, we consider the $\varepsilon=0$ limit of system (1), (2), which corresponds to its $(1+\sum n_{j})$-dimensional fast subsystem (1), with $I_{\mathrm{ext}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$, where $V_{\mathrm{h}}$ is now treated as a parameter. For fixed $V_{\mathrm{h}}$ and voltages in the range $-80\ldots 30$ mV of interest, system (1), (2) has only stable steady states (no limit cycles, that is, no neuronal spikes); see SI for details. For a fixed hold voltage $V_{\mathrm{h}}$, the steady states of (1), (2) with $I_{\mathrm{ext}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$ satisfy the algebraic equations for $(V,x_{j})_{j=1,\ldots,N}$
$\begin{split}\textstyle{\sum_{j}}I_{j}(x_{j},V)&=g_{\mathrm{c}}(V_{\mathrm{h}}-V),\\ x_{j}&=x_{\infty,j}(V),\end{split}$ (4)
where $I_{\mathrm{eq}}(V)=\sum_{j}I_{j}(x_{\infty,j},V)$ is the equilibrium current for fixed membrane potential $V$. The solutions of (4) form a (1D) steady-state curve of system (1), (2), which is normally hyperbolic (transversally attracting) for $\varepsilon=0$. For $\varepsilon\neq 0$ the increase of $V_{\mathrm{h}}$ with speed $\varepsilon\Delta_{V}$ introduces a slow variation of all states $(V,x_{j})$. Hence, $V$ and the feedback current $I_{\mathrm{vc}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$ (as measured), are not at their steady-state values given by (4), but they are changing dynamically. This results in a difference between the measured curve $(I_{\mathrm{vc}},V_{\mathrm{h}})$ in Fig. 1(b) and the $(I,V)$-values of the desired steady-state curve (4).
Figure 3: VC and CC in silico. Protocol as described for Fig. 1 applied to (a) a type-I Morris-Lecar neuron model [30] as example of excitatory cell, and (b) a type-I Wang-Buzsaki neuron model [31] as example of inhibitory interneuron; see SI for parameter values. The two-dimensional fast subsystem has a S-shaped steady-state curve satisfying (4). The steady-state curve (4) and the (multi-color) S-shaped curve from the VC protocol indistinguishable throughout the range of input currents $I_{\mathrm{vc}}$. The orange curve resulting from the CC protocol is very close to $S^{0}$ and the VC protocol near its dynamically stable parts.
Figure 4: VC and CC protocols applied to the Morris-Lecar model in a parameter regime where it behaves as a type-II neuron model.
Geometrical singular perturbation theory (GSPT) by Fenichel [32] implies that after decay of initial transients every trajectory of the VC protocol modelled by system (1), (2) satisfies eq. (4) for the steady-state curve up to order $\varepsilon$. The first-order terms in $\varepsilon$ are
$\begin{split}I_{\mathrm{eq}}(V)-I_{\mathrm{vc}}(V)&\approx I_{\mathrm{eq}}^{\prime}(V)\dot{V}\,\frac{\tau_{1/2}}{\ln 2}\\ &\lesssim I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})0.2\mbox{\,mV},\end{split}$ (5)
where $\tau_{1/2}$ is the time for deviations from the transversally stable steady-state curve (4) to decay to half of their initial value; see SI for details (eq. (S7) and Fig. S3). We estimate $\tau_{1/2}$ from recovery transients after disturbances naturally occuring from imperfections in the voltage clamp during VC runs as $\tau_{1/2}\lesssim 0.075$ s (see SI, Fig. S3(b)), such that $\dot{V}\tau_{1/2}/\ln 2\approx\dot{V}_{\mathrm{h}}\tau_{1/2}/\ln 2\lesssim 0.2$ mV. Thus, the systematic bias between $I_{\mathrm{vc,sm}}(V_{\mathrm{h}})$ in Fig. 1(b) and the true steady state curve $I_{\mathrm{eq}}(V)$ caused by dynamically changing $V_{\mathrm{h}}$ is below measurement disturbances.
Estimate (5) can also be tested in silico. Figures 3 and 4 emulate both VC and CC protocols with the Morris-Lecar model [30], a biophysical model typically used for excitatory neurons of either type-I (Fig. 3(a)) or type-II (Fig. 4) excitability, and with the Wang-Buzsaki model (Fig. 3(b)), which is a classical model of inhibitory interneuron. Both models are of form (1) ($j=n_{j}=1$). For the chosen parameter set (see SI), the curves $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$ and $(I_{\mathrm{eq}}(V_{\mathrm{h}}(t)),V_{\mathrm{h}}(t))$ are order $\varepsilon$ ($\approx 1$%) apart.
Consequently, time profile $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$ follows closely the curve (4) of stable steady states of the fast subsystem (1) with VC protocol $I_{\mathrm{ext}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$, treating $V_{\mathrm{h}}$ as a parameter. We now connect the curve (4) to a curve of fast-subsystem equilibria of the CC protocol, which is in part unstable. To this end we recast the VC protocol in the form of a CC protocol with non-constant current ramp speed $\varepsilon\Delta_{I,\mathrm{vc}}(t)$ and disturbances: the smoothed time profile $I_{\mathrm{vc,sm}}(V_{\mathrm{h}}(t))$ in Fig. 1(b) of the VC run (thick, in blue/brown/red) equals the raw-data measured time profile $I_{\mathrm{vc}}(t)$ (thin blue curve with fluctuations) plus disturbances $\eta_{\mathrm{vc}}(t)$, defined by $\eta_{\mathrm{vc}}(t)=I_{\mathrm{vc,sm}}(V_{\mathrm{h}}(t))-I_{\mathrm{vc}}(t)$. After smoothing, the derivative $I_{\mathrm{vc,sm}}^{\prime}(V_{\mathrm{h}})$ w.r.t. $V_{\mathrm{h}}$ is moderate ($\lesssim 20$ pA///mV in modulus at its maximum near $I_{\mathrm{db}}$), such that the VC protocol implies
$I_{\mathrm{ext}}\approx I_{\mathrm{vc,sm}}+\eta_{\mathrm{vc}},\ \dot{I}_{\mathrm{vc,sm}}=\varepsilon\Delta_{I,\mathrm{vc}}(t),$ (6)
where $\Delta_{I,\mathrm{vc}}(t)=I_{\mathrm{vc,sm}}^{\prime}(V_{\mathrm{h}}(t))\Delta_{V}$, hence $\Delta_{I,\mathrm{vc}}(t)\approx I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}}(t))\Delta_{V}$, with upper bound $\max_{t}|\Delta_{I,\mathrm{vc}}(t)|\lesssim 3.7$ pA///ms in the range of Fig. 1(b). Thus, $I_{\mathrm{vc,sm}}$ is indeed still slow. Hence, except for disturbances $\eta_{\mathrm{vc}}(t)$, the external current $I_{\mathrm{ext}}$ is slowly varying according to a CC protocol with slowly time-varying speed $\varepsilon\Delta_{I,\mathrm{vc}}(t)$, such that the VC protocol (2) is equivalent to the CC protocol (6) with disturbances $\eta_{\mathrm{vc}}$.
Systems (1), (3) and (1), (6) without disturbances ($\eta_{\mathrm{vc}}=0$) are both models of CC protocols. They have the same fast subsystem (1) when setting $\varepsilon=0$ (i.e., for constant input current) and identifying $I_{\mathrm{h}}$ and $I_{\mathrm{vc,sm}}$. The respective fast-subsystem steady states $(V,x_{j})_{j=1,\ldots,N}$ satisfy
$\begin{split}\textstyle{\sum_{j}}I_{j}(x_{j},V)&=I_{\mathrm{h}}(=I_{\mathrm{vc,sm}}),\\ x_{j}&=x_{\infty,j}(V).\end{split}$ (7)
However, systems (1), (3) and (1), (6) differ by the nature of their respective slow variables $I_{\mathrm{h}}$ and $I_{\mathrm{vc,sm}}$: $I_{\mathrm{h}}$ is an externally applied hold current for (3), while $I_{\mathrm{vc,sm}}$ is a measured (and smoothed) current from the feedback control $g_{\mathrm{c}}(V_{\mathrm{h}}-V)$ of the voltage source for (6). Thus, while the S-shaped steady-state curve $(I_{\mathrm{eq}}(V),V)$ is identical for both systems, it contains large unstable segments as a steady-state curve of (1), (3), while it is always stable as a steady-state curve of (1), (6). The change in stability is caused by the disturbances $\eta_{\mathrm{vc}}$, which are current adjustments generated by the feedback term in (2), $I_{\mathrm{vc}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$. Along most of the curve $(I_{\mathrm{vc,sm}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ the $\eta_{\mathrm{vc}}$ are small fluctuations such that $I_{\mathrm{vc,sm}}(V_{\mathrm{h}})\approx I_{\mathrm{vc}}(V_{\mathrm{h}})$ and the feedback is approximately non-invasive [7]. Estimate (5) ensures that the measurements $I_{\mathrm{vc}}(V)$ stay close to $I_{\mathrm{eq}}(V)$. Therefore, we can conclude that the VC protocol (2) with slowly varying feedback reference signal $V_{\mathrm{h}}$ reveals the entire family of steady states of a neuron (type 1 or 2) with constant external current $I_{\mathrm{ext}}$, both stable (observable) and unstable (non-observable, hidden). Consequently, the N-shaped $I$-$V$ relations for type-1 neurons reported in [19, 20, 23] equal S-shaped steady-state bifurcation diagrams for these neurons with respect to $I_{\mathrm{ext}}$. They are tractable with a VC protocol where the current $I_{\mathrm{ext}}$ is a sufficiently slowly varying feedback current with sufficiently small fluctuations $\eta_{\mathrm{vc}}$. In particular, this allows us to detect and pass through fold bifurcations directly in the experiment.
Figure 5: Periodic spiking responses to step-current protocols for PY cells 5 and 1. Panels (a1-b1) are reproductions of Fig. 1(a) of the main manuscript and of Fig. S1(a), respectively, with a dashed rectangle chosen around $I_{\mathrm{h}}=200$ pA. The corresponding zoomed view is provided in panels (a2-b2). Panels (a3-b3) show the time profiles of voltage responses $V_{\mathrm{cc}}$ from a current-clamp stimulation with constant hold current $I_{\mathrm{h}}=200$ pA; the red dashed lines highlight the transient convergence towards a periodic response, which we use to compute $\tau_{1/2}$. Finally, panels (a4-b4) show a phase-plane projection of the time series shown in panels (a3-b3), namely plotting $V_{\mathrm{cc}}^{\prime}$ against $V_{\mathrm{cc}}$. This shows that the responses stabilise onto a periodic solution.
In contrast, for the CC open-loop protocol (3) applying a slowly varying electrical current the neuron dynamically transitions between its observable rest states and its observable (dynamically stable) spiking states (see Fig. 1(b) orange time profile). The speed of variation $\varepsilon\Delta_{I}\lesssim 0.75\times 10^{-3}$ pA///ms is such that $I_{\mathrm{ext}}$ varies by only about $1$ pA per spiking period. Transients to step current responses in the stable spiking region have a half-time for decay $\tau_{1/2}\lesssim 0.2$ s (see SI, Fig. S5 for the response to a step to $I_{\mathrm{ext}}=200$ pA), permitting us to estimate the systematic bias between true steady-state spiking and the response to dynamically changing $I_{\mathrm{ext}}$, similar to (5) for certain features of the periodic spike train, as described below.
As the description of the protocol in the main manuscript states we stimulate the neuron under current clamp with a current step during a calibration phase. This is done for a number of values of the step current. Panels (a2) and (b2) in Fig. 5 show the time profiles of the recorded voltage $V_{\mathrm{cc}}$ for a current step to $I_{\mathrm{h}}=200$ pA for PY cells 5 and 1. Panels (a1) and (a2) show where the end point of the step ($I_{\mathrm{h}}=200$ pA) is located in the respective bifurcation diagrams (vertical dashed line) for PY cells 5 and 1, respectively. The two dots along the vertical dashed line at $I_{\mathrm{h}}=200$ pA in panels (a1) and (a2) mark the maximum and minimum values of the voltage response (after transient) to the step-current stimulation, as shown in panels (a3) and (b3); panels (a2) and (b2) are zoomed views of panels (a1) and (b1), respectively, near $I_{\mathrm{h}}=200$, where the current-clamp data with constant hold current are considered. We observe that both responses to this step-current stimulation lead to a periodic spiking behaviour, after initial transients have decayed. This is especially visible on the bottom plots, obtained by approximating the time derivative of the voltage time series, which is plotted against the voltage, hence giving a phase diagram. In both case, the phase-plane trajectory reveals a convergence towards a periodic (spiking) state (inner oscillations). The transient (outer oscillations) seems to indicate a weak stability of these spiking states; a blue line indicates the direction of time, with two dots corresponding to points at times $t_{n}$ and $t_{n+m}$.
The black dots in panel (a1) indicate that for PY cell 5 there is quantitative agreement between the response to the CC protocol with slowly-varying applied current $I_{\mathrm{h}}(t)$ and the response to the current step. There is less agreement for PY cell 1, between step response and the response to the CC protocol with slow variation of $I_{\mathrm{h}}$. The time profile suggests that there are either different possible spiking responses and, hence, more bifurcations in the dynamic protocol, such as period-doubling bifurcations, or the condition of the PY cell has drifted between step response calibration and CC ramp.
We use the decay rate of transients toward the stable spiking response in panels (a3) and (b3) to estimate time to half-decay of transients $\tau_{1/2}$ for spiking responses in the stable part of the spiking region (similar to the estimate for $\tau_{1/2}$ from the large-fluctuation decay in Fig. S3 in Section S3 for eq. (5)).
Observing how the spike maxima settle to their steady-state value in Fig. 5(a3), we conservatively estimate the half-time for the decay to half the amplitude deviation from steady-state maximum as $\tau_{1/2}\leq 0.2$ s. This half-time enters estimates for the bias caused by varying the hold current $I_{\mathrm{h}}$ dynamically with speed $\varepsilon\Delta_{I}$. For example, the estimate for the bias in the minimum voltage $V_{\min}$ of the spike equals to first order in $\varepsilon$ is
$\operatorname{bias}(V_{\min})\approx V_{\min}^{\prime}(I_{\mathrm{h}})\times\frac{\mathrm{d}}{\mathrm{d}t}I_{\mathrm{h}}(t)\times\frac{\tau_{1/2}}{\ln 2}.$
We can see in Fig. 5(a) that $V_{\min}(I_{\mathrm{h}})$ changes in the spiking region with about $8$ mV per $120$ pA, so $V_{\mathrm{min}}^{\prime}(I)\approx 0.07$ mV/\!//pA. The current changes with $\mathrm{d}I/\mathrm{d}t\approx 7$ pA///s. Thus, to first order in $\varepsilon$, the error then amounts to:
$\operatorname{bias}(V_{\min})\approx V_{\min}^{\prime}(I_{\mathrm{h}})\times\frac{\mathrm{d}}{\mathrm{d}t}I_{\mathrm{h}}(t)\times\frac{\tau_{1/2}}{\ln 2}\approx 0.14\mbox{\,mV.}$.
Hence, the effect from changing the current $I_{\mathrm{h}}$ dynamically is small. The bias estimate grows to infinity when the amplitude $V_{\max}(I_{\mathrm{h}})$ envelope assumes a square-root like shape and $\tau_{1/2}$ goes to infinity, as is the case at the Hopf bifurcation.
Thus, combining the VC protocol (6) for varying $I_{\mathrm{vc,sm}}$, and the CC protocol (3) for varying $I_{\mathrm{h}}$, enables us to interpret the data sets from both protocols in Fig. 1(b) as a bifurcation diagram including unstable states.
For the experimental curves presented in Fig. 1(b,c) (and also in Fig. S1), the disturbances $\eta_{\mathrm{vc}}$ are not small in some unstable parts of the reported steady-state curve (e.g., near $V_{\mathrm{h}}=-30$ mV in Fig. 1(b)), caused by imperfect voltage clamping across the membrane (see Fig. SI3(b) in SI for a zoom into these intermittent dropout events). Furthermore, the distances between $(I_{\mathrm{vc,sm}}(t),V_{\mathrm{h}}(t))$ in the VC run and $(I_{\mathrm{h}}(t),V_{\mathrm{cc}}(t))$ in the CC run for the same cell are visibly larger along parts of the curve corresponding to dynamically stable stationary states. This is due to the natural drift of the neuron’s physiological properties as it changes conductance properties (e.g., temperature, degradation,..) over time.
Conclusion.–
Tracking non-observable (hidden) states and their stability boundaries in experimental settings bridges the gap between real-world phenomena and nonlinear science. Specifically, closed-loop control methods with slow variations of feedback reference signals enable to dissect the underlying states of multi-scale complex systems. Following this strategy, we demonstrate the feasibility of tracking a family of neuronal steady states (stable and unstable) via variations of reference signal and reparameterizing the obtained curve by feedback current. Crucially, the method is versatile in that it applies to different cell types, both pyramidal (PY) neurons (like in Fig. 1), of excitability type I or II (Figs. 3(a) and 4), and inhibitory interneurons (Fig. 3(b)).
These results can be extended to more complex neuronal states (e.g., limit cycles) by combining two recent advances, namely: 1. dynamic-clamp electrophysiology [33, 34, 35, 2], which allows for a two-way real-time communication between a neuronal tissue (e.g., neuron, network or neuronal sub-processes) and a computer simulation of a component of the neuronal tissue. This introduces the full range of feedback control theory and engineering techniques into electrophysiology; 2. Control-based continuation for experiments (CBCE, which combines feedback control theory and pseudo-arclength numerical continuation for tracking solution branches of nonlinear systems directly from noisy experimental data [7]. Indeed, CBCE method has been successfully applied in various experimental systems, for instance, in mechanical vibration and buckling experiments [8], pedestrian flow experiments [9], cylindrical pipe flow simulations [11] and feasibility studies for synthetic gene networks [12, 13]. Noteworthy, the methodology presented in this work can be seen as a special case of CBCE. Future work will focus on employing the full CBCE to track more complex neuronal states. This is crucial as it will allow to validate computational models by comparing their numerical bifurcation diagrams with experimental ones. This approach will have direct impact in experimental labs enabling biologists to have access to novel states with which to carry out new experimental paradigms. We envisage this extended protocol could be used to develop next-generation closed-loop deep-brain stimulation devices to treat certain brain diseases such as epilepsy [36], where it could help monitor control the excitability threshold of key neural populations in real time.
Materials and Method.–
Protocols.–
Male Wistar rats were used in this study (age P21, N=8 animals). For the bifurcation diagrams (see Fig. 1(b) and Fig. SI1), 4 of these 8 rats were used. PY Cells 1-3 were recorded from different animals, PY cells 4-5 were recorded from the same animal. To study the effects of hysteresis (see Fig. 2) 20 neurons were recorded from the 4 remaining animals (221222, 230110, 230111, 230112). The control group were 6 neurons from 2 rats (230110, 230112). The QX group were 8 neurons from 4 rats (221222, 230110, 230111, 230112). The QX+Cd group were 6 neurons from 4 rats (221222, 230110, 230111, 230112).
The use and handling of animals was performed in accordance with the European Community Council Directive 86/609/EEC. Horizontal 300-µm-thick brain slices were prepared as described in our previous studies [37]. The slices contained the hippocampus and the adjacent cortical regions and were kept in the artificial cerebrospinal fluid (ACSF) with the following composition (in mM): 126 NaCl, 24 NaHCO3, 2.5 KCl, 2 CaCl2, 1.25 NaH2PO4, 1 MgSO4, 10 glucose (bubbled with 95% O2/5% CO2 gas mixture). All the listed chemicals were purchased from Sigma-Aldrich (St. Louis, MO, USA).
We performed the whole-cell patch-clamp recordings of the principal neurons in the entorhinal cortex. Neurons within the slices were visualized using a Zeiss Axioscop 2 microscope (Zeiss, Oberkochen, Germany), equipped with a digital camera (Grasshopper 3 GS3-U3-23S6M-C; FLIR Integrated Imaging Solutions Inc., Wilsonville, OR, USA) and differential interference contrast optics.
[…]
We applied first VC and then CC protocol to $5$ neurons in the entorhinal cortex of $4$ male Wistar rats [37]. We first performed the VC neuronal recordings varying the hold voltage $V_{\mathrm{h}}$ from $-80$ mV to $+30$ mV slowly with $\dot{V}_{\mathrm{h}}=1.83$mV///s, while measuring current, called $I_{\mathrm{vc}}$ in Fig. 1(b,c). Subsequently, for the CC recordings we first determine the minimal injected current required to induce a depolarization block of action potential generation ($I_{\mathrm{db}}$ in Fig. 1(b), upper limit of current input where firing occurs). Then we injected hold current $I_{\mathrm{h}}$, gradually increasing from 00 pA to $I_{\mathrm{db}}$ during $60$ s, such that $\dot{I}_{\mathrm{h}}$ is in the range $6.6\ldots 7.4$ pA///s for the $5$ neurons, while recording voltage $V_{\mathrm{cc}}(t)$. See SI for further checks (e.g., for hysteresis).
Computational models.–
We applied the VC and the CC protocols, with slow variations in the feedback reference signal and in the applied current, respectively, to the classical neuron model adapted to many contexts, namely the Morris-Lecar model [30] (ML), and to a standard neuron model that was specifically designed to account for inhibitory neural activity, namely the Wang-Buzsaki model [31] (WB).
The ML equations are as follows
$\begin{split}C\dot{V}&=-g_{\mathrm{L}}(V-V_{\mathrm{L}})-g_{Ca}m_{\infty}(V)(V-V_{Ca})-g_{K}w(V-V_{K})+g_{c}(V_{\mathrm{h}}-V),\\ \dot{w}&=\phi\frac{w_{\infty}(V)-w}{\tau_{w}(V)},\end{split}$ (8)
with the following voltage-dependent (in)activation and time-constant functions:
$\begin{split}m_{\infty}(V)&=0.5(1+\tanh((V-V_{1})/V_{2})),\\ w_{\infty}(V)&=0.5(1+\tanh((V-V_{3})/V_{4})),\\ \tau_{w}(V)&=\cosh((V-V_{3})/(2V_{4}))^{-1}.\end{split}$ (9)
To obtain Figs. 3(a) and 4, we have used the following parameter values.
Table 1: Parameter values for the Morris-Lecar model (8).
parameter $C$ $g_{L}$ $V_{L}$ $g_{Ca}$ $V_{Ca}$ $g_{K}$ $V_{K}$ $g_{c}$ $\phi$ $V_{1}$ $V_{2}$ $V_{3}$ $V_{4}$
unit pF nS mV nS mV nS mV nS mV mV mV mV
value type-I 20 2 -60 4.0 120 12 -84 40 0.067 -1.2 18 12 17.4
value type-II 20 2 -60 4.4 120 12 -84 150 0.040 -1.2 18 2 30.0
Finally, for both the VC protocol with slow variation, and the CC protocol with slow variation, the speed of the variation was chosen to be equal to $\varepsilon=0.01$. Note that $g_{c}$ can be decreased to 7, which is in the same order of magnitude as the experimental one. Moreover, the capacitance chosen for the model simulations is on the same order of magnitude as observed in the experiments. Nevertheless, the key point to note is that we are not aiming for quantitative agreement since this is a conductance-based phenomenological model.
The WB equations are as follows
$\begin{split}C\dot{V}&=-g_{\mathrm{L}}(V-V_{\mathrm{L}})-g_{Na}m_{\infty}^{3}(V)h(V-V_{Na})-g_{K}n^{4}(V-V_{K})+g_{c}(V_{\mathrm{h}}-V),\\ \dot{h}&=\phi\frac{h_{\infty}(V)-h}{\tau_{h}(V)},\\ \dot{n}&=\phi\frac{n_{\infty}(V)-n}{\tau_{n}(V)},\end{split}$ (10)
with the following voltage-dependent (in)activation and time-constant functions:
$\begin{split}\alpha_{m}(V)&=0.1(V+35)/(1-\exp(-0.1(V+35)))\\ \beta_{m}(V)&=4.0\exp(-0.0556(V+60))\\ \alpha_{h}(V)&=0.07\exp(-0.05(p+58))\\ \beta_{h}(V)&=1/(1+\exp(-0.1(V+28)))\\ \alpha_{n}(V)&=0.01(V+34)/(1-\exp(-0.1(V+34)))\\ \beta_{n}(V)&=0.125\exp(-0.0125(V+44))\\ p_{\infty}(V)&=\alpha_{p}(V)/(\alpha_{p}(V)+\beta_{p}(V)),\;p\in\{m,h,n\},\\ \tau_{p}(V)&=1/(\alpha_{p}(V)+\beta_{p}(V)),\;p\in\{m,h,n\}\end{split}$ (11)
To obtain Fig. 3(b) of the main manuscript, we have used the following parameter values.
Table 2: Parameter values for the Wang-Buzsaki model (10).
parameter $C$ $g_{L}$ $V_{L}$ $g_{Na}$ $V_{Na}$ $g_{K}$ $V_{K}$ $g_{c}$ $\phi$
unit pF nS mV nS mV nS mV nS
value 1 0.1 -65 35.0 55 9 -90 20 5
Finally, for both the VC protocol with slow variation, and the CC protocol with slow variation, the speed of the variation was chosen to be equal to $\varepsilon=0.01$.
Data and code availability.–
All presented data, as well as, codes to process them, are available via the following Github repository: https://github.com/serafimrodrigues/Amakhin-etal-2025/.
S1 Text.
Supplementary text and Supplementary Figures.
Acknowledgments.–
SR acknowledges the grant PID2023-146683OB-100 funded by
MICIU/AEI /10.13039/501100011033 and by ERDF, EU. Additionally, SR acknowledges support from Ikerbasque Foundation and the Basque Government through the BERC 2022-2025 program and by the Ministry of Science and Innovation: BCAM Severo Ochoa accreditation CEX2021-001142-S / MICIU / AEI / 10.13039/501100011033. Moreover, SR acknowledges the financial support received from BCAM-IKUR, funded by the Basque Government by the IKUR Strategy and by the European Union NextGenerationEU/PRTR, as well as, support of SILICON BURMUIN no. KK-2023/00090 funded by the Basque Government through ELKARTEK Programme.
Supplementary Information (SI)
S1 Experimental bifurcation diagrams for Pyramidal (PY) cells 1–4, including current and voltage clamp ramp data
Figure S1: Experimental bifurcation diagrams for PY cells 1–4. Protocols identical to Fig. 1(b) in main manuscript.
Figure S1 presents the recordings for VC and CC protocol for PY cells 1–4 in the same way as Fig. 1(b) in the main manuscript (which was for PY cell 5). The graphs superimpose the observations of our VC protocol with slow variation of the feedback reference signal and the CC protocol with slow variation of the externally applied currents. The graphs confirm (as does Fig. 1(c) from the main manuscript) that type-II properties of the neurons vary between individual cells. We also observe that in all PY cells for $I_{\mathrm{h}}>I_{\mathrm{db}}$ and for PY cell 3 for small $I_{\mathrm{h}}$ the distance between the stable steady states obtained from the VC protocol and the corresponding trajectory segments obtained from the CC protocol is larger than the bias caused by dynamic variation of $V_{\mathrm{h}}$ and $I_{\mathrm{h}}$ respectively, based on estimate (5) of the main text. We do not have a complete physiological explanation for this mismatch, but we hypothesise that this is due to the natural physiological drift of underlying neuronal processes as one performs successive recordings on the same PY cell (see also Fig. 2 of the main manuscript for a comparison between successive ramp-up and ramp-down on the same PY cell).
S2 Spiking frequency and interspike intervals during CC protocol
A characteristic for the class of excitability is the dependence of the spiking frequency on the external current $I_{\mathrm{ext}}$ as $I_{\mathrm{ext}}$ approaches the lower limit of the spiking region, with a drop in frequency expected for type-I neurons. Fig. S2 presents the interspike interval as a function of the hold current $I_{\mathrm{h}}$ in panel (a), and the frequency against $I_{\mathrm{h}}$ in panel (b) for the 5 PY cells for which experimental bifurcation diagrams are shown in Fig. 1 of the main manuscript and Fig. S1. We detected crossing times for $V_{\mathrm{cc}}(t)$ through the threshold $V=0$ from below (or equivalently Poincaré map, at $V=0$). As visible in Fig. S2, even though the 5 PY cells have different steady-state curves (obtained with the VC protocol), their frequency behaviours during the CC protocol are qualitatively similar with drops from $\approx 15$ Hz to $\lesssim 5$ Hz for all PY cells. Hence, the frequency drop does not allow us to conclusively distinguish between different excitability classes for the present PY cells. However, the steady-state bifurcation diagrams obtained through VC protocols show that PY cells 3-5 exhibit an experimental bifurcation diagram that is compatible with type-1 excitability, while PY cells 1 and 2 have type-2 compatible steady-state curves.
Figure S2: Interspike intervals in seconds (panel (a)) and frequency in Hz (panel (b)) for the spiking responses during the CC protocols with slowly-varying applied current near the low-current limit of the spiking region.
S3 Details of intermittent fluctuations around unstable steady states
When performing the VC protocol with slow variation of the feedback reference signal we observe fluctuations around the states that would be unstable in the CC protocol. The highlighted ellipse in Fig. S3(a) shows a prominent example of large-scale fluctuations. We observe this phenomenon in all PY cells. Figure S3(b) shows details of the typical time profile of these fluctuations in the measured feedback current $I_{\mathrm{vc}}=g_{c}(V_{h}-V)$ in blue. In the background Fig. S3(b) shows a neuronal firing pattern (in orange) around the same currents to give a visual comparison of fluctuation amplitude and frequency with the natural neuronal firing patterns. We also compute the moving average of $I_{\mathrm{vc}}(t)$ (dashed blue), which shows the equilibrium level of $I_{\mathrm{vc}}$. We observe that between the large fluctuations the feedback current $I_{\mathrm{vc}}$ stabilises to this equilibrium level with a consistent rate of decay and consistently smaller levels of smaller fluctuations. We used the observed decay rate toward the mean to estimate the decay half-time $\tau_{1/2}$ in eq. (5) of the main manuscript.
To simulate the intermittent fluctuations around unstable steady states observed in the experimental recordings (see Fig. S3), we used the Morris-Lecar model (eq. (8) of the main manuscript) with added external current spikes in the $V$-equation. Namely, we replaced the $V$-equation from eq. (8) by
$\begin{split}C\mathrm{d}V&=\bigg{(}-g_{\mathrm{L}}(V-V_{\mathrm{L}})-g_{Ca}m_{\infty}(V)(V-V_{Ca})-g_{K}w(V-V_{K})\ldots\\ &\ldots+g_{c}(V_{\mathrm{h}}-V)+\sum_{i}I_{i}(t)\bigg{)}\mathrm{d}t+\sigma_{W}\mathrm{d}W\end{split}$ (S1)
where $I_{i}(t)=I_{p}\delta(t-t_{i})$ is a current spike modelled by a Dirac delta input, which we approximated by two Heaviside functions. Namely: $\delta(t-t_{i})\approx\text{Heav}(t-t_{i})\ast\text{Heav}(t_{i}+t_{\mathrm{dur}}-t)$, with the duration of the spike chosen to be $t_{\mathrm{dur}}=10$ms, and its amplitude being $I_{p}=50$pA. We also added noise in the form of a Wiener process $W$ of amplitude $\sigma_{W}=0.5$. Other parameters are as in Table 1. For the loss of contact simulation (Fig. S4(b)) we used a piecewise-constant $g_{c}$ dropping from $20$ to $5$ along the unstable branch of equilibria (other parameters as in Table 1) and we also added noise using a Wiener process $W$ as in the previous case, of amplitude $\sigma_{W}=0.1$.
Figure S3: Experimental intermittent fluctuations. (a) Reproduction of Fig. 1(b) of main manuscript with the region of intermittent fluctuations highlighted by a black ellipse. (b) Fluctuation of the current $I_{\mathrm{vc}}$ from the VC protocol, with apparent intermittent fluctuation (blue), compared with the voltage spikes obtained from the CC protocol (orange). Red dashed lines illustrate how we compute $\tau_{1/2}$.
Figure S4: Reproducing the experimental intermittent fluctuations with the ML model. (a-b) Simulation of the ML model with current spikes artificially added to the voltage equation, viewed in the $(I_{\mathrm{vc}},V_{\mathrm{h}})$-plane in panel (a) and with the time series of $I_{\mathrm{vc}}$ in panel (b). (c) Simulation of the ML model with artificial loss of contact through a drop of gain $g_{\mathrm{c}}$, viewed in the $(I_{\mathrm{vc}},V_{\mathrm{h}})$-plane.
We hypothesise that these intermittent fluctuations are currents (possibly mediated by voltage-gated sodium and/or calcium channels) generated in distal neuronal sites (e.g., dendrites, axons) due to non-ideal voltage clamp (i.e., space-clamp error due to failure to achieve isopotential across the entire neuronal cell).
This can be understood by first comparing I-V relationships of voltage-gated channels of neurons in slice preparations versus isolated neurons, and subsequently inferring what occurs to isolated voltage-gated channels that are poorly controlled due to space-clamp error. Neurons in slice preparations have lower input resistance and higher membrane capacitance when compared to an isolated neuron [38]. We observed the same effect with our entorhinal cortical neurons, which in slice preparations had $R\in\left]150,350\right[M\Omega$ and $C\approx 100$ pF, while in a previous study (unpublished) the isolated case had $R\in\left]500,1000\right[M\Omega$ and $C\in\left]10,30\right[$ pF. Thus, in slice preparations the typically observed voltage range in which the I-V relationship has a negative slope (e.g., for sodium channels) is quite large [39]. In ideal conditions (i.e., when voltage-gated channels are more or less clustered within a neuronal site, for instance, close to the soma), space-clamp error is minimised. In this case, when the membrane voltage is slowly increased then individual negative peak currents of voltage-gated channels are successively activated (i.e., a kind of averaging effect occurs due to the clustering of channels), which give rise to a wide voltage range in which the I-V relationship has a negative slope. In other words, the current peaks are distributed across the voltage range of the I-V relationship [39].
Noteworthy, the fast sodium current peaks occur during this negative slope phase because the voltage dependence of the transient sodium channel activation (which mostly mediate these peaks) is similar to the activation of the persistent sodium channels. The poorly voltage-clamped transient sodium channels mediate the fast negative currents, while the transient ones mediate a slow negative current, both of which appear in the same voltage range [40]. In contrast, isolated neurons typically lose most of their thin dendritic branches (i.e., abbreviated morphology), hence voltage-gated channels (e.g., sodium current) display only one negative peak threshold current under a step depolarisation [38, 41]. From these observations, it is reasonable to assume that in non-ideal conditions (i.e., space-clamp error) there are isolated channels furthest from the cell soma or electrode contact (e.g., distal dendrites with thin branches, or dendritic spines with increased resistance of the cytoplasm, or distal sodium channels) [42, 43]. These isolated channels have I-V relationships with negative slope but with a shorter voltage range and distinct activation threshold. Thus, under a VC protocol the mismatch between the hold voltage ($V_{\mathrm{h}}$) and the real-membrane voltage ($V$) of distal voltage-gated channels can be very large, resulting in the observed intermittent fluctuations. Additionally to the above mentioned factors, the delay of the feedback (associated with the amplifier, which is in the order tens of nanoseconds) is likely to play a role in these observed intermittent fluctuations. However, from a theoretical standpoint, a fixed delay can induce a Hopf bifurcation, and therefore, robust oscillations. So instead, we envisage a varying delay, due to possibly latent processes, could create this effect. Of course, to really address this point, further experiments are required.
To simulate the above hypothesis we added additional instantaneous currents (i.e., to model the effect of non-ideally clamped channels as external forcing that correspond to non-homogeneous terms); see Section Computation models of the main manuscript for details, in particular eq. (10). The results are depicted in Fig. S4 in the $(I,V)$-plane (panel (a) and as a corresponding time series $I_{\mathrm{vc}}(t)$ (panel (b)). Both, time profile and $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$-curve closely resemble the fluctuations in the experimental data in Fig. S3(a). Although, these simulations may explain the experimental observations, we do not exclude other possible mechanisms (or a mix of mechanisms). To this end, note that the standard VC protocol was designed to stabilise the membrane voltage $V$ at some static reference signal $V_{h}$, for a given, large enough, control gain $g_{c}$. However, our VC protocol involves slowly varying $V_{h}$, which causes the neuronal membrane $V$ to quasi-statically stabilise (drift) without ever reaching the true steady state. Along the unstable branch, the repulsion (associated with unstable eigenvalues) may not necessarily be constant and thus the degree of effectiveness of the control signal $g_{c}(V_{h}-V)$ may vary. Noteworthy, this control signal is effectively the proportional control term in the proportional–integral–derivative (PID) control formalism. Thus, it is also possible the proportional control does not achieve a full control of more complex physiological processes in which case the control may only be intermittent. Other possibilities may include temporary loss of contact between the electrode and the neuronal membrane. To elucidate these points, we simulated equation (8) but with variations in $g_{c}$. Specifically, we considered a piecewise-constant $g_{c}$ corresponding to decreasing its default value (thus modelling loss of control) along the unstable branch of equilibrium; see Fig. S4(c) for a simulation. While this scenario bears less resemblance to the experimental observation, it can either be used to rule out this possibility or to consider new experiments to further clarify these observations.
S4 Experiments estimating hysteresis effects and blocking of voltage-gated sodium and calcium ion channels
The voltage-clamp recordings of the responses to the slow ramps of command voltage (Fig. S5 (a)) were made to obtain the current-voltage relationships (I-V relationships). With the potassium gluconate-based pipette solution being used, the I-V relationship obtained by applying the ascending ramp was not the same as the one obtained by applying the descending ramp (Fig. S5 (b)). The ascending ramp produced an S-shaped I-V relationship, which had the negative slope in the voltage range of $-50$ to $-30$ mV (Fig. S5 (b), CTRL, green trace). The descending ramp of the command voltage produced an I-V curve with a positive slope throughout the physiological voltage range.
Figure S5: (a) The representative voltage-clamp recordings of the responses to the slow ramps of the command potential. Top traces represent the whole-cell current, bottom trace represent the command potential protocol. Dark green lines correspond to the ascending ramp; light green lines correspond to the descending ramp. (b) Pairs of averaged ascending and descending I-V relationships, obtained in control conditions (left panel), with QX314 in the pipette solution (middle panel), and with both QX314 in the pipette solution and Cd2+ in the ACSF (right panel). Gray bars represent the standard error. The I-V relationships were normalized to the value of current at 0 mV. C. The comparison of I-V slope ratios in three experimental conditions (Kruskal-Wallis ANOVA, P$<0.001$, followed by Dunn’t post hoc test; - P$<0.05$, ∗∗ - P$<0.01$)
We hypothesized that the mismatch between the ascending and descending I-V relationships can be attributed to the activity of the slow inactivating sodium and calcium ion channels. These slowly inactivating sodium and calcium channels are active at the ascending ramp, whereas high depolarization at the descending ramp inactivates them such that they are ineffective and thus do not contribute to the S-shaped I-V relationship. It has previously been reported that persistent sodium channels are present in deep entorhinal cortex neurons, and that current through these channels can be detected even when the rate of membrane depolarization is very low [44]. A prepulse to depolarized voltages inactivates persistent sodium channels of the neocortical neurons [40]. T-Type calcium channels are also known to produce the “window” currents due to the overlap of the activation and steady-state inactivation curves [45]. In order to check this hypothesis we used QX314 (6 mM) and Cd2+ (50 µM), which are the blockers of the voltage-gated sodium and calcium ion channels, respectively. To quantify the effects of the applied substances on the mismatch between the ascending and descending I-V relationships we introduced the I-V slope ratio (Fig. S5 (c)). It was calculated as the ratio of the minimal slope of the ascending I-V relationship to the slope of the descending I-V relationship at -35 mV. In control conditions, the I-V slope ratio was negative due to a strong mismatch between the I-V relationships (Fig. S5 (b,c), CTRL). When the pipette solution containing QX314 was used, the observed mismatch between the I-V relationships was much less pronounced (Fig. S5 (b), middle panel), resulting in the significantly higher I-V slope ratio. With QX314 in the pipette solution and Cd2+ in the ACSF, both the ascending and descending I-V relationships were monotonically increasing functions (Fig. S5 (b), right panel), resulting in positive values of the I-V slope ratios. The results obtained suggest that the negative slope of the ascending I-V relationship is mainly mediated by persistent sodium channels and to a lesser extent by voltage-gated calcium channels.
In the model the channel properties depend sensitively on the maximal conductances (e.g., $g_{\mathrm{Ca}}m_{\infty}$ for calcium in eq. (8) of the main manuscript). A small change in these parameters in eq. (8) removes the two fold points from the steady-state curve (in a so-called cusp bifurcation), resulting in a change from type-I to type-II behavior.
S5 Equilibrium analysis of the general conductance-based model for the $\mathbf{V}$-clamped protocol
For a fixed hold voltage $V_{\mathrm{h}}$ a general conductance-based model, written using indices $j$ for ${n_{\mathrm{c}}}$ channels and rates $r_{j}(V)=1/\tau_{j}(V)$, reads (see main text, equation (1)):
$C\dot{V}$ $=-\sum_{j=1}^{n_{\mathrm{c}}}I_{j}(x_{j},V)+g_{\mathrm{c}}(V_{\mathrm{h}}-V),$ $\dot{x}_{j}$ $=(x_{\infty,j}(V)-x_{j})r_{j}(V).$ (S2)
This section states a sufficient criterion for the uniqueness and linear stability of equilibria for system (S2). Equilibria are at $x_{\mathrm{eq},j}=x_{\infty,j}(V_{\mathrm{eq}})$, where $V_{\mathrm{eq}}$ satisfies
$V_{\mathrm{eq}}$ $=V_{\mathrm{h}}-\frac{1}{g_{\mathrm{c}}}\sum_{j=1}^{n_{\mathrm{c}}}I_{j}(x_{\infty,j}(V_{\mathrm{eq}}),V_{\mathrm{eq}})=:\operatorname{alg}_{\operatorname{fix}}(V_{\mathrm{eq}}).$ (S3)
We denote the ${n_{\mathrm{c}}}\times 1$ vectors ($j=1,\ldots,{n_{\mathrm{c}}}$)
$I_{\infty,j}$ $=I_{j}(x_{\infty,j}(V_{\mathrm{eq}}),V_{\mathrm{eq}}),$ $x_{\infty,j}^{\prime}$ $=x_{\infty,j}^{\prime}(V_{\mathrm{eq}}),$
$I_{x,j}$ $=\partial_{1}I_{j}(x_{\infty,j}(V_{\mathrm{eq}}),V_{\mathrm{eq}}),$ $r_{\mathrm{eq},j}$ $=r_{j}(V_{\mathrm{eq}}),$
$I_{V,j}$ $=\partial_{2}I_{j}(x_{\infty,j}(V_{\mathrm{eq}}),V_{\mathrm{eq}}),$ $r_{\mathrm{eq},j}^{\prime}$ $=r_{j}^{\prime}(V_{\mathrm{eq}}),$
$\tau_{\mathrm{eq},j}$ $=\tau_{j}(V_{\mathrm{eq}}).$
The derivative of $\operatorname{alg}_{\operatorname{fix}}(V)$ in $V=V_{\mathrm{eq}}$ is (using the notation $\bm{1}=(1,\ldots,1)^{\mathsf{T}}$)
$\operatorname{alg}_{\operatorname{fix}}(V_{\mathrm{eq}})^{\prime}$ $=-\frac{1}{g_{\mathrm{c}}}[I_{x}^{\mathsf{T}}x_{\infty}^{\prime}+I_{V}^{\mathsf{T}}\bm{1}]=:\mu_{\mathrm{c}},$
which is dimensionless. Consequently, a sufficient condition for the uniqueness of the fixed point $V_{\mathrm{eq}}$ is $|\mu_{\mathrm{c}}|<1$, or, which is guaranteed by the estimate
$m_{\mathrm{c}}:=\frac{1}{g_{\mathrm{c}}}[\|I_{x}.x_{\infty}^{\prime}\|_{1}+\|I_{V}\|_{1}]<1,$
where (.)(.)( . ) means element-wise multiplication (so, $v.w=\operatorname{diag}(v)w$), and $\|v\|_{1}:=|v_{1}|+\ldots+|v_{n_{\mathrm{c}}}|$ is the $1$-norm of a vector $v$. The condition $m_{\mathrm{c}}<1$ is also sufficient for linear stability of $(V_{\mathrm{eq}},x_{\mathrm{eq},(\cdot)})$: the Jacobian of (S2) in $(V_{\mathrm{eq}},x_{\mathrm{eq},(\cdot)})$ is
$J_{\mathrm{eq}}:=\begin{bmatrix}-(g_{\mathrm{c}}+I_{V}^{\mathsf{T}}\bm{1})/C,&-I_{x}^{\mathsf{T}}/C\\[2.15277pt] r_{\mathrm{eq}}.x_{\infty}^{\prime},&-\operatorname{diag}(r_{\mathrm{eq}})\end{bmatrix},$
(note that the matrix entries are in row order $1\times 1$, $1\times{n_{\mathrm{c}}}$, ${n_{\mathrm{c}}}\times 1$ and ${n_{\mathrm{c}}}\times{n_{\mathrm{c}}}$, and that the term $\operatorname{diag}(r_{\mathrm{eq}}^{\prime})x_{\infty}(V_{\mathrm{eq}})$ cancels in the bottom left entry). By rescaling the $x_{j}$ to $x_{j}/(x_{\infty,j}^{\prime}(1+\delta)$ with an arbitrary $\delta>0$ we can make the lower ${n_{\mathrm{c}}}$ rows diagonally dominant. This gives then in the upper right entry $-(I_{x}.x_{\infty}^{\prime})^{\mathsf{T}}(1+\delta)/C$:
$J_{\mathrm{scal}}:=\begin{bmatrix}-(g_{\mathrm{c}}+I_{V}^{\mathsf{T}}\bm{1})/C,&-\frac{1+\delta}{C}(I_{x}.x_{\infty}^{\prime})^{\mathsf{T}}\\[2.15277pt] r_{\mathrm{eq}}/(1+\delta),&-\operatorname{diag}(r_{\mathrm{eq}})\end{bmatrix},$
Typically, the derivative of $I_{j}$ with respect to $V$ is non-negative (so, $I_{V,j}\geq 0$). Hence, the matrix satisfies Gershgorin’s theorem for some $\delta>0$ if
$g_{\mathrm{c}}+\sum_{j=1}^{n_{\mathrm{c}}}I_{V,j}$ $>\sum_{j=1}^{n_{\mathrm{c}}}|I_{x,j}x_{\infty,j}^{\prime}|\mbox{,\quad or\quad}g_{\mathrm{c}}+\|I_{V}\|_{1}>\|I_{x}.x_{\infty}^{\prime}\|_{1}.$
This is equivalent to $m_{\mathrm{c}}<1$. The fixed-point relation (S3) also ensures that
$V_{\mathrm{h}}(1-m_{\mathrm{c}})\leq V_{\mathrm{eq}}\leq\frac{V_{\mathrm{h}}}{1-m_{\mathrm{c}}},$
which gives a bound on the relationship between $V_{\mathrm{h}}$ and $V_{\mathrm{eq}}$, that becomes sharper for larger $g_{\mathrm{c}}$. Hence, for conductance-based models with clamped voltage $V_{\mathrm{h}}$ the difference between the clamp voltage $V_{\mathrm{h}}$ and the unique stable equilibrium potential across the membrane is of order $m_{\mathrm{c}}$ (so gets small for large gain $g_{\mathrm{c}}$).
S6 Conditions for validity of estimate for systematic bias due to dynamic ramp of voltage
Estimate (5) in the main text provides an approximation to first order in $\varepsilon$ for the bias between steady-state current, $I_{\mathrm{eq}}(V_{\mathrm{h}})$, for fixed hold voltage $V_{\mathrm{h}}$, and the actually measured current $I_{\mathrm{vc}}(t)$ when the hold voltage is changed slowly with $\dot{V}_{\mathrm{h}}=\varepsilon\Delta_{V}$:
$I_{\mathrm{eq}}(V_{\mathrm{h}}(t))-I_{\mathrm{vc}}(t)\approx I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\dot{V}_{\mathrm{h}}/\alpha.$ (S4)
In this expression $\alpha$ is the decay rate toward the equilibrium $I_{\mathrm{eq}}(V_{\mathrm{h}})$ at fixed $V_{\mathrm{h}}$ and $I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})$ is the slope of the equilibrium (steady-state) curve $I_{\mathrm{eq}}(V_{\mathrm{h}})$ we are interested in. In practice, we estimate
$I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}}(t))$ $\approx I_{\mathrm{vc}}^{\prime}(V_{\mathrm{h}}(t)),$ $\alpha^{-1}$ $\approx\tau_{1/2}/\ln 2,$ (S5)
where $I_{\mathrm{vc}}(V_{\mathrm{h}})$ is the measured $(I,V)$-curve during voltage clamp runs (VC) and $\tau_{1/2}$ is the time of decay to half of the transients during large fluctuations of measured $I_{\mathrm{vc}}$ during VC runs (see Section S3 and Fig. S4). The estimates (S5) are themselves accurate to first order in $\varepsilon$, such that their bias is only a second-order effect in (S4) if $\alpha>0$. Pretending $I_{\mathrm{vc}}(t)$ was governed by a scalar first-order ordinary differential equation $\dot{I}_{\mathrm{vc}}=f_{IV}(I_{\mathrm{vc}},V_{\mathrm{h}})$, bias estimate (S4) would follow from implicit differentiation with respect to time and $I_{\mathrm{vc}}$ , where $\alpha=-\partial_{1}f_{IV}(I_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})>0$.
Justifying (S4) based on a more accurate model for the VC experiment such as a higher-dimensional conductance-based model requires that two conditions are met:
  • presence of a dominant eigenvalue with spectral gap in the linearization near the steady-state curve, and
  • observability of fluctuations with the slowest rate of decay in the measured current.
Examples of such models are equations (1), (2) in the main text; see also equation (8) for the concrete Morris-Lecar model used for Fig. 2 in the main text and Fig. S4. More precisely, let us collect the entire vector of conductance-based model variables into $y=(V,x_{j})_{j=1,\ldots,n_{\mathrm{c}}}$, such that the conductance-based model (1), (2) for voltage clamp with a slow ramp of hold voltage $V_{\mathrm{h}}$ from the main text is of the general form
$\dot{y}$ $=f(y,V_{\mathrm{h}}),$ $\dot{V}_{\mathrm{h}}$ $=\varepsilon\Delta_{V}.$ (S6)
For each fixed hold voltage $V_{\mathrm{h}}$ (so, at $\varepsilon=0$) the voltage clamp stabilizes an equilibrium $y_{\mathrm{eq}}(V_{\mathrm{h}})$, such that $0=f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$.
Spectral gap
The first condition for validity of (S4) is that among the eigenvalues of the matrix $J(V_{\mathrm{h}}):=\partial_{1}f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ (which all have negative part), there is one dominant eigenvalue $-\alpha(V_{\mathrm{h}})$ of $J(V_{\mathrm{h}})$ with largest real part. This dominance is expressed as a spectral gap condition for $J(V_{\mathrm{h}})$, which means that all other eigenvalues must have real part less than some bound $-\beta(V_{\mathrm{h}})<-\alpha(V_{\mathrm{h}})<0$ with the condition
$\beta^{-1}(V_{\mathrm{h}})$ $\ll\alpha^{-1}(V_{\mathrm{h}})\sim\mbox{\ order 1.}$. (S7)
The current flowing through the membrane is the scalar output $I_{\mathrm{vc}}(y)=\sum_{j}I_{j}(x_{j},V)\in\mathbb{R}$, depending on the state variables of the conductance-based model, which we measure. Near the equilibrium $y_{\mathrm{eq}}(V_{\mathrm{h}})$ this output is to first order of the form
$I_{\mathrm{vc}}(y)\approx I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))+\partial I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))[y-y_{\mathrm{eq}}(V_{\mathrm{h}})]\mbox{,\quad where $\partial I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))=:c(V_{\mathrm{h}})^{\mathsf{T}}\in\mathbb{R}^{1\times(n_{\mathrm{c}}+1)}$ is a row vector.}$.
Observability
The second condition for validity of (S4) is that the right eigenvector $y_{\mathrm{ev}}(V_{\mathrm{h}})$ for the dominant eigenvalue $-\alpha(V_{\mathrm{h}})$ has a non-zero (more precisely, non-small) projection in variations of this measured output:
$\partial I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))y_{\mathrm{ev}}(V_{\mathrm{h}})=c(V_{\mathrm{h}})^{\mathsf{T}}y_{\mathrm{ev}}(V_{\mathrm{h}})\neq 0\mbox{.}$ (S8)
If the dominant direction $y_{\mathrm{ev}}$ is observable according to (S8), then we may normalize the eigenvector $y_{\mathrm{ev}}$ such that $c(V_{\mathrm{h}})^{\mathsf{T}}y_{\mathrm{ev}}(V_{\mathrm{h}})=1$ for all $V_{\mathrm{h}}$. We denote the left eigenvector of $J(V_{\mathrm{h}})$ for the dominant eigenvalue $-\alpha(V_{\mathrm{h}})$ by $y_{\mathrm{ad}}(V_{\mathrm{h}})$, normalized such that $y_{\mathrm{ad}}(V_{\mathrm{h}})^{\mathsf{T}}y_{\mathrm{ev}}(V_{\mathrm{h}})=1$. If $J(V_{\mathrm{h}})$ satisfies the spectral gap condition (S7), its inverse is $J(V_{\mathrm{h}})^{-1}\approx-\alpha(V_{\mathrm{h}})y_{\mathrm{ev}}(V_{\mathrm{h}})y_{\mathrm{ev}}(V_{\mathrm{h}})^{\mathsf{T}}$ (ignoring small terms of order $\beta^{-1}$), thus, it is close to the rank-$1$ matrix $-\alpha y_{\mathrm{ev}}y_{\mathrm{ad}}^{\mathsf{T}}$. As $y_{\mathrm{eq}}(V_{\mathrm{h}})$ is the steady-state curve of (S6) (so, $0=f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$), we have (abbreviating $\partial_{2}f_{\mathrm{eq}}(V_{\mathrm{h}}):=\partial_{2}f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ and dropping argument $V_{\mathrm{h}}$ from $\alpha$, $c$, $y_{\mathrm{ev}}$, $y_{\mathrm{ad}}$, $J$, and $\partial_{2}f_{\mathrm{eq}}$)
[…]
$I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})$ $=\partial I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))y_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})=c^{\mathsf{T}}y_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\approx-\alpha y_{\mathrm{ad}}^{\mathsf{T}}\partial_{2}f_{\mathrm{eq}}\mbox{,\quad implying}$
$y_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})$ $\approx y_{\mathrm{ev}}I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})$
when neglecting terms of order $\beta^{-1}$. Expansion of (S6) near the steady-state curve $(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ implies that solutions $y(t)$ of (S6) satisfy (after transient have decayed, dropping the argument $t$ from $V_{\mathrm{h}}$)
$y(t)-y_{\mathrm{eq}}(V_{\mathrm{h}}(t))\approx J^{-1}y_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\dot{V}_{\mathrm{h}}\approx-\alpha y_{\mathrm{ev}}y_{\mathrm{ad}}^{\mathsf{T}}y_{\mathrm{ev}}I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\dot{V}_{\mathrm{h}}=-\alpha y_{\mathrm{ev}}I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\dot{V}_{\mathrm{h}},$
ignoring terms of order $\varepsilon^{2}$ and $\varepsilon\beta^{-1}$. Inserting this expansion into the expansion for the output $I_{\mathrm{vc}}(y(t))$ gives the desired first-order bias estimate,
$I_{\mathrm{vc}}(y(t))-I_{\mathrm{eq}}(V_{\mathrm{h}}(t))\approx\partial I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))[y(t)-y_{\mathrm{eq}}(V_{\mathrm{h}}(t))]\approx-\alpha c^{\mathsf{T}}y_{\mathrm{ev}}I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\dot{V}_{\mathrm{h}}=-\alpha I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\dot{V}_{\mathrm{h}}.$ (S9)
Finally, we observe that the measured current $I_{\mathrm{vc}}(t)$ when $V_{\mathrm{h}}$ changes dynamically is not $-\sum_{j}I_{j}(x_{j},V)$ but (as measured at the electrode) equal to $g_{\mathrm{c}}(V_{\mathrm{h}}-V)$. However, $I_{\mathrm{vc}}(t)=g_{\mathrm{c}}(V_{\mathrm{h}}(t)-V(t))=\sum_{j}I_{j}(x_{j},V)+C\dot{V}(t)$, where $\dot{V}$ is a component of $\dot{y}\approx f(y,V_{\mathrm{h}})\approx J[y(t)-y_{\mathrm{eq}}(V_{\mathrm{h}}(t))]$, which equals $y_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\dot{V}_{\mathrm{h}}$ and is of order $\varepsilon$ after transients have decayed. Thus, the correction due to this difference in (S9) is only another $O(\varepsilon)$ term in $I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})$, making it a second-order effect overall.